Consider a game between an employer (E) and a worker (W). The worker can choose to work hard (H) or slack (S); the employer can dock the worker's pay (D) or not (N). The worker gets enjoyment worth a 1 from slacking, but hurt worth − 2 from docked pay. Thus, a worker who works hard and whose pay is not docked gets a 0; one who slacks and whose pay is docked gets 1 − 2 = −1; and so on. Morale will be so low in the office that docking pay comes at a cost for the employer. The employer gets −2 from the worker slacking and −1 from docking pay.

  1. Set up this game as a simultaneous-move game, and find the equilibrium.
  2. Next, suppose that the worker chooses H or S first and that the employer chooses D or N after having observed the worker's action. Draw the game tree and find the subgame-perfect Nash equilibrium.
  3. Now suppose that before the worker acts, the employer can commit to a strategy. For example, the threat "D if S" ("If you slack, I will dock your pay"). How many such strategies does the employer have? (Make sure each strategy is complete. In other words, the threat "D if S" is actually the strategy - DN - D if you slack, N if you don't slack.) Write the table for this game. Find all pure-strategy Nash equilibria.
  4. How do your answers to parts (b) and (c) differ? Use what you know about game theory to explain the difference.
Answer #1

a) It expresses

  • Itexpresses the initial concentration, the change in concentration, and the equilibrium concentration for each species in the reaction. From the chart you can determine the changes in the concentrations of each species and the equilibrium concentrations. From the example, we start with the folowing information.

NO

H2

N2

H2O

Initial Concentration (M)

0.100

0.0500

0

0.100

Change in Concentration (M) -2x -2x +x +2x
Equilibrium Concentration (M)

0.062

There are three types of equilibrium: stable, unstable, and neutral . In economics, economic equilibrium is a situation in which economic forces such as supply and demand are balanced and in the absence of external influences the (equilibrium) values of economic variables will not change.

b) Here if workers chooses H or F then it is H-S =2-1=1

Nash equilibrium is a concept within game theory where the optimal outcome of a game is where there is no incentive to deviate from their initial strategy.Overall, an individual can receive no incremental benefit from changing actions, assuming other players remain constant in their strategies.Nash equilibrium also allows for the possibility that decision makers follow randomised strategies. Allowing for randomisation is important for the mathematics of game theory because it guarantees that every (finite) game has a Nash equilibrium.

c) Here the employer do have a different strategy for his own employees and strategy DN-D=2*1-1=21-1=1

Six elements of a successful employer

1.Get a clear understanding of your existing employer brand first.

2.Create clear values and a company mission – and communicate it.

3.Define EVPs for different candidate/employee groups.

4.Gain external recognition for your efforts.

5.Tell your employees' stories.

6.Encourage social media advocacy.

Think of all of the value your employees gain from working for you. This might be compensation packages, flexible work time, opportunities for advancement, cool office place perks, or an intellectually stimulating environment.

Employer branding is the firm's corporate image or culture created to attract and retain the type of employees the firm is seeking. Through employer branding, people get to know what the company stands for, the people it hires, the fit between jobs and people, and the results it recognizes and rewards.

d) Game theory is the  theory  that deals with the determination of the optimal strategies for each of the players in a game with well-defined rules. In a typical scenario, two players, A and B, are playing a game. Player A is required to make a decision in ignorance of a simultaneous decision made by player B. The outcome is a consequence of the two decisions. In a zero-sum game A wins what B loses. Many real-life situations can be modelled in game-theoretic terms. The rules used by the players to determine their strategies are called decision rules. Game Theory. Game theory is the study of the ways in which interacting choices of economic agents produce outcomes with respect to the preferences (or utilities) of those agents, where the outcomes in question might have been intended by none of the agents.